Find the maximum or minimum vaule of the function and the vaule of x when it occurs.
y= 4x^2+2x-1

Answer by Alan3354(69443)

Find the maximum or minimum vaule of the function and the vaule of x when it occurs.
y= 4x^2+2x-1
------------
It's on the line of symmetry: x = -b/2a = -2/8
x = -1/4
y = -5/4
Minimum at (-1/4,-5/4)